# Power to the Clients: Federated Learning in a Dictatorship Setting

### v1

@@ Appendix B Algorithm 3: Cheater Client
Appendix C What if the dictator client does not have the learning rate?
@@
Appendix D Proof of Algorithm 2
@@
Using Eq. 13, we can write $\theta_{2}$ as the following:
@@
Using Eq. 16 we can write $\theta_{t+1}$ as the following:
@@
As it can be seen in Eq. 19, the exact final weights with our method would be $\hat{\theta}^{\mathcal{P}}_{T+1}-\eta{\sum\nolimits_{n=1,n\notin{\mathcal{P}}}^{N}\nabla{\mathcal{L}_{n}(\theta_{T})}}$ where $\hat{\theta}^{\mathcal{P}}_{T+1}$ would represent the weights for the final iteration if only the $P$ collaborative dictator clients were contributing to the system during the training procedure and like the other clients never existed. Again, the term $\eta{\sum\nolimits_{n=1,n\notin{\mathcal{P}}}^{N}\nabla{\mathcal{L}_{n}(\theta_{t})}}$ is negligible since it is the updates only for one iteration and can not affect the final model too much, especially if the model achieved by the collaborative dictators is robust to such perturbations. Moreover, because of the nature of FL, the dictator clients are always one step behind and can not cancel this residual term. However, one could come up with more sophisticated methods in order to approximate or predict this residual term by observing the gradients through the training process.
Appendix E Proof of Algorithm 3
@@
$=\hat{\theta}^{\mathcal{P}}_{E-1}-\eta(\sum\nolimits_{k\in\mathcal{P}}\nabla{\mathcal{L}_{k}(\hat{\theta}^{\mathcal{P}}_{E-1})}+{\sum\nolimits_{n=1,n\notin\mathcal{P}}^{N}\nabla{\mathcal{L}_{n}(\theta_{E-1})}})$
$\quad-\eta(\sum\nolimits_{i=1}^{E-1}\nabla{\mathcal{L}_{1}(\hat{\theta}^{1}_{i})}-\sum\nolimits_{i=1}^{E-1}(\nabla{\mathcal{L}_{1}(\hat{\theta}^{\mathcal{P}}_{t})}+\nabla{\mathcal{L}_{2}(\hat{\theta}^{\mathcal{P}}_{t})})+M^{2}_{t}+{\sum\nolimits_{n=1,n\notin\mathcal{P}}^{N}\nabla{\mathcal{L}_{n}(\theta_{E})}}).$
Using Eq. 21 we will have:
@@
Finally, from Eq. 22 we have:
$\theta_{E+1}$ $=\hat{\theta}^{1}_{E-1}-\eta(\sum\nolimits_{k\in\mathcal{P}}\nabla{\mathcal{L}_{k}(\hat{\theta}_{E-1}^{\mathcal{P}})}+{\sum\nolimits_{n=1,n\notin\mathcal{P}}^{N}\nabla{\mathcal{L}_{n}(\theta_{E-1})}})$
@@
Appendix F Experiments for NLP
For NLP experiments, we used the distilbert-base-uncased Sanh et al. (2019) model for text classification as the global model and selected the AG news datasetZhang et al. (2015) which has has 4 different labels. Hence, we considered a FL with four clients for this case where each client has samples of only one label.
F.1 Single Dictator Client
Figure 7 demonstrates the loss function of global model when there is no dictator client and when client 1 becomes a dictator. Table 3 demonstrates accuracy of the global model when each client becomes dictator. We can see that each client has successfully dominated the training and led the global model to learn only that client’s dataset.
@@ F.1 Single Dictator Client
Table 3: Performance of the global model on each local dataset for AG news dataset and the single dictator client scenario.
F.2 Collaborative Dictator Clients
Figure 8 demonstrates the loss function of global model when two or three clients become collaborative dictators. Table 4 demonstrates accuracy of the global model for these cases. We can see that the collaborative dictators successfully dominated the training and led the global model to learn only their dataset.
@@ F.2 Collaborative Dictator Clients
Table 4: Performance of the global model on each local dataset for AG news dataset and the collaborative dictator clients scenario.
Appendix G Main results with 1-sigma error bars
In this section, we present extended versions of our main results with additional statistical details. All experiments were repeated using 5 different random seeds to account for variability. Tables 5 and 6 provide expanded versions of Table 1, reporting the mean and the corresponding 1-sigma error bars. Similarly, Tables 7 and 8 offer detailed results corresponding to Table 2.
@@ Appendix G Main results with 1-sigma error bars
Table 5: Performance of the global model on each local dataset for MNIST and the single dictator client scenario.
@@
Table 6: Performance of the global model on each local dataset for CIFAR10 and the single dictator client scenario.
Method [0,1] [2,3] [4,5] [6,7] [8,9]
Table 7: Performance of the global model on each local dataset for MNIST and the collaborative dictator clients scenario.
Method [0,1] [2,3] [4,5] [6,7] [8,9]
@@
Table 8: Performance of the global model on each local dataset for CIFAR10 and the collaborative dictator clients scenario.

### v2

@@ G-A Experiments for Betrayal in Collaboration
Appendix H What if the dictator client does not have the learning rate?
@@
Appendix I Proof of Algorithm 2
@@
Using Eq. (14), we can write $\theta_{2}$ as the following:
@@
Using Eq. (17) we can write $\theta_{t+1}$ as the following:
@@
As it can be seen in Eq. (20), the exact final weights with our method would be $\hat{\theta}^{\mathcal{P}}_{T+1}-\eta{\sum\nolimits_{n=1,n\notin{\mathcal{P}}}^{N}\nabla{\mathcal{L}_{n}(\theta_{T})}}$ where $\hat{\theta}^{\mathcal{P}}_{T+1}$ would represent the weights for the final iteration if only the $P$ collaborative dictator clients were contributing to the system during the training procedure and like the other clients never existed. Again, the term $\eta{\sum\nolimits_{n=1,n\notin{\mathcal{P}}}^{N}\nabla{\mathcal{L}_{n}(\theta_{t})}}$ is negligible since it is the updates only for one iteration and can not affect the final model too much, especially if the model achieved by the collaborative dictators is robust to such perturbations. Moreover, because of the nature of FL, the dictator clients are always one step behind and can not cancel this residual term. However, one could come up with more sophisticated methods in order to approximate or predict this residual term by observing the gradients through the training process.
Appendix J Proof of Algorithm 3
@@
$=\hat{\theta}^{\mathcal{P}}_{E-1}-\eta(\sum\nolimits_{k\in\mathcal{P}}\nabla{\mathcal{L}_{k}(\hat{\theta}^{\mathcal{P}}_{E-1})}+{\sum\nolimits_{n=1,n\notin\mathcal{P}}^{N}\nabla{\mathcal{L}_{n}(\theta_{E-1})}})-\eta(\sum\nolimits_{i=1}^{E-1}\nabla{\mathcal{L}_{1}(\hat{\theta}^{1}_{i})}-\sum\nolimits_{i=1}^{E-1}(\nabla{\mathcal{L}_{1}(\hat{\theta}^{\mathcal{P}}_{t})}$
$\quad+\nabla{\mathcal{L}_{2}(\hat{\theta}^{\mathcal{P}}_{t})})+M^{2}_{t}+{\sum\nolimits_{n=1,n\notin\mathcal{P}}^{N}\nabla{\mathcal{L}_{n}(\theta_{E})}}).$
Using Eq. (22) we will have:
@@
Finally, from Eq. (23) we have:
$\theta_{E+1}$ $=\hat{\theta}^{1}_{E-1}-\eta(\sum\nolimits_{k\in\mathcal{P}}\nabla{\mathcal{L}_{k}(\hat{\theta}_{E-1}^{\mathcal{P}})}+{\sum\nolimits_{n=1,n\notin\mathcal{P}}^{N}\nabla{\mathcal{L}_{n}(\theta_{E-1})}})-\eta(M^{2}_{t}+{\sum\nolimits_{n=1,n\notin\mathcal{P}}^{N}\nabla{\mathcal{L}_{n}(\theta_{E})}}).$
@@
Appendix K Experiments for NLP
For NLP experiments, we used the distilbert-base-uncased [21] model for text classification as the global model and selected the AG news dataset[30] which has has 4 different labels. Hence, we considered a FL with four clients for this case where each client has samples of only one label.
K-A Single Dictator Client
Figure 7 demonstrates the loss function of global model when there is no dictator client and when client 1 becomes a dictator. Table II demonstrates accuracy of the global model when each client becomes dictator. We can see that each client has successfully dominated the training and led the global model to learn only that client’s dataset.
@@
TABLE II: Performance of the global model on each local dataset for AG news dataset and the single dictator client scenario.
K-B Collaborative Dictator Clients
Figure 8 demonstrates the loss function of global model when two or three clients become collaborative dictators. Table III demonstrates accuracy of the global model for these cases. We can see that the collaborative dictators successfully dominated the training and led the global model to learn only their dataset.
@@
TABLE III: Performance of the global model on each local dataset for AG news dataset and the collaborative dictator clients scenario.
Appendix L Results for random client dropping
Method MNIST
[0,1] [2,3] [4,5] [6,7] [8,9]
Dictator client: 11 99.62 0.00 0.00 0.00 0.00
Dictator client: 22 0.00 89.76 0.00 0.00 0.00
Dictator client: 33 0.00 0.00 96.53 0.00 0.00
Dictator client: 44 0.00 0.00 0.00 98.84 0.00
Dictator client: 55 0.00 0.00 0.00 0.00 94.35
Dictator clients: 22,33 0.00 83.84 81.75 0.00 0.00
TABLE IV: Performance of the global model on each local dataset for the MNIST dataset under single dictator client and collaborative dictator clients scenarios when one client is randomly dropped during each update.
Figure 9: Loss function on each client’s dataset for one dictator client for the random client dropping experiment.
We conducted experiments in both the single-dictator and collaborative-dictator settings under a more realistic scenario in which a randomly selected client is dropped in each training round. As shown in Table IV, the dictator client (or collaborative dictators) continues to consistently override benign client contributions, and the attack remains effective despite client dropout. The primary observable effect of this randomness is a slightly noisier training trajectory, with minor oscillations in the loss curve induced by the stochastic removal of clients (Figure 9).
Appendix M Dictator clients against gradient norm clipping Defense
Method MNIST CIFAR-10
[0,1] [2,3] [4,5] [6,7] [8,9] [0,1] [2,3] [4,5] [6,7] [8,9]
Dictator clients: $2,3$ 0.00 88.19 87.80 0.00 0.00 0.00 35.08 43.17 0.00 0.00
Dictator clients: $2,3$ + norm clipping 0.43 65.40 51.28 10.62 0.00 13.00 17.25 23.50 0.85 24.65
Dictator clients: $2,3,4$ 0.00 84.87 80.22 94.19 0.00 0.00 18.38 40.02 46.05 0.00
Dictator clients: $2,3,4$ + norm clipping 0.00 69.78 52.35 87.01 0.00 0.00 27.90 22.25 37.85 1.55
TABLE V: Performance of the global model on each local dataset for MNIST and CIFAR-10 under collaborative dictator-client attacks, with and without server-side gradient norm clipping as defense.
We evaluate our attack under server-side gradient norm clipping, a common defense in federated learning that limits the size of each client’s update before aggregation to reduce the impact of abnormal or malicious gradients. In the single-dictator setting, this defense is effective and stops the attack by shrinking the dictator client’s update. However, in the collaborative-dictator setting, where multiple dictator clients participate together, the defense becomes less effective. As the number of dictator clients increases and they form a majority, their clipped updates still combine to overpower the benign clients, allowing the attack to remain successful. The results are shown in Table V.
Appendix N Compute Resources
We conducted all experiments using a single NVIDIA H100 GPU. Reproducing the main results requires 65 GB of VRAM, while the NLP experiments used up to 75 GB.
Appendix O Main results with 1-sigma error bars
In this section, we present extended versions of our main results with additional statistical details. All experiments were repeated using 5 different random seeds to account for variability. Tables VI and VII provide expanded versions of Table I, reporting the mean and the corresponding 1-sigma error bars.
@@
TABLE VI: Performance of the global model on each local dataset for MNIST and the single dictator client and collaborative dictator clients scenarios.
@@
TABLE VII: Performance of the global model on each local dataset for CIFAR10 and the single dictator client and collaborative dictator clients scenarios.